The table below lists the prices from last year and the base year for a college-related basket of goods. Assume that the typical basket of goods for a college student consists of 210 gallons of gas, 55 pizzas, 40 6-packs, and 5 textbooks.

Basket of College-Related Goods Basket of Goods Price Base Year (dollars) Price Last Year (dollars) Gasoline (per gallon) $ 1.60 $ 2.05 Pizza (per pizza) 4.05 7.80 Beer (per 6-pack) 4.05 6.82 Textbook (per book) 97.00 227.00

Instructions: Round your answers to two decimal places. a. Using the values above, what is the rate of inflation between the base year and last year? %

b. Assume that rather than buying textbooks for their courses last year, all students decided to buy online access cards at $95 per textbook. What is the rate of inflation between the base year and last year now? %
Answer #1

(a)

Cost of basket, last year = 210 x 1.6 + 55 x 4.05 + 40 x 4.05 + 5 x 97 = 336 + 222.75 + 162 + 485 = 1205.75

Cost of basket, this year = 210 x 2.05 + 55 x 7.8 + 40 x 6.82 + 5 x 227 = 430.5 + 429 + 272.8 + 1135 = 2267.3

CPI last year = 100 (base year CPI is always 100)

CPI this year = (Cost of basket, this year / Cost of basket, last year) x 100 = (2267.3/1205.75) x 100 = 188.04

Inflation rate = % Change in CPI = (188.04/100) - 1 = 1.8804 - 1 = 0.8804 = 88.04%

(b)

Cost of basket, last year = 210 x 1.6 + 55 x 4.05 + 40 x 4.05 + 5 x 95 = 336 + 222.75 + 162 + 475 = 1195.75

Cost of basket, this year = 210 x 2.05 + 55 x 7.8 + 40 x 6.82 + 5 x 95 = 430.5 + 429 + 272.8 + 475 = 1607.3

CPI last year = 100 (base year CPI is always 100)

CPI this year = (Cost of basket, this year / Cost of basket, last year) x 100 = (1607.3/1195.75) x 100 = 134.42

Inflation rate = % Change in CPI = (134.42/100) - 1 = 1.3442 - 1 = 0.3442 = 34.42%